# ViBiDSampler: Enhancing Video Interpolation Using Bidirectional Diffusion Sampler

### v2

@@ 3.1 Bidirectional Sampling
This approach effectively constrains the sampling process for bounded generation between the start frame () and the end frame (). As depicted in Fig. 3 (b), our method seamlessly connects the temporally forward and backward paths so that the sampling trajectory stays within the SVD manifold, resulting in smooth and coherent transitions throughout the interpolation process.
@@ 3.2 Additional manifold guidances
Last frame guidance with DDS. DDS (Chung et al., 2023) synergistically combines the diffusion sampling and Krylov subspace methods (Liesen & Strakos, 2013) such as the conjugate gradient (CG) method, guaranteeing the on-manifold solution of the following optimization problem:
@@
Here, we leverage the DDS framework to guide a start-frame-conditioned sampling path of SVD to a start- and end-frame-conditioned sampling path. Specifically, for the temporally forward path, conditioned on the start frame (), we take the DDS step on denoised estimate to guide the last frame of to align with . For the temporally backward path, conditioned on the end frame (), we take the DDS step on denoised estimate to guide the last frame of to align with . In practice, we set as last frame extractor, as matched condition which are for temporally forward path and for temporally backward path:
@@
While SVD replaces text embedding with CLIP image embedding, we can still use CFG++ for image-to-video diffusion models to ensure better image-video alignment. Specifically, after applying CFG++ into SVD sampling framework, the Euler-step of SVD (3) now reads:
where the last term in (3) is replaced by . In practice, we apply DDS guidance before CFG++ update, so in (13) should be replaced with as in (12). We experimentally found that incorporating DDS and CFG++ guidance improves the interpolation performance of bidirectional sampling. The overall sampling method effectively steers the SVD sampling path to perform keyframe interpolation in an on-manifold manner, fully leveraging the generation capabilities of SVD. The detailed algorithm is provided in Algorithm 1. The vanilla bidirectional sampling can be implemented by removing DDS guidance (orange) and replacing the CFG++ update (blue) with a traditional CFG update. The detailed algorithm of the vanilla bidirectional sampling is provided in Appendix A.
@@ 4.2 Comparative studies
Quantitative evaluation. For quantitative evaluation, we used LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017) to assess the quality of the generated frames, and FVD (Unterthiner et al., 2019) to evaluate the overall quality of the generated videos. As shown in Table 1, our method surpasses the other baselines in terms of fidelity. Moreover, it achieves superior perceptual quality, particularly in scenarios involving dynamic motions (DAVIS), indicating that our approach effectively addresses the issue of deviations from the diffusion manifold, resulting in improved video generation quality.
@@ Appendix A Algorithm
Appendix B Additional experimental results

### v3

@@ 3.1 Bidirectional Sampling
Note that the amount of renoising is determined by the noise difference between and , which is crucial to avoid mode collapsing behavior from the forward path sampling. This approach effectively constrains the sampling process for bounded generation between the start frame () and the end frame (). As depicted in Fig. 3 (b), our method seamlessly connects the temporally forward and backward paths so that the sampling trajectory stays within the SVD manifold, resulting in smooth and coherent transitions throughout the interpolation process.
@@ 3.2 Additional manifold guidances
Last frame guidance with DDS. We introduce additional frame guidance to achieve more accurate bounded generation. Specifically, we employ DDS guidance to align the last frame with in the temporally forward path and with in the temporally backward path. DDS (Chung et al., 2023) synergistically combines the diffusion sampling and Krylov subspace methods (Liesen & Strakos, 2013) such as the conjugate gradient (CG) method, guaranteeing the on-manifold solution of the following optimization problem:
@@
Here, we present a novel insight that DDS, originally designed to address image inverse problems, can also be effectively applied to bounded video generation. Specifically, we define as a last-frame extractor and as the target condition, which corresponds to for the temporal forward path and for the temporal backward path in bounded video generation. Then we take DDS step from (11) on denoised estimates ( and ) with target condition ( and ) to get the on-manifold solutions ( and ) of the following optimization problems:
@@
While SVD replaces text embeddings with CLIP image embeddings, we empirically found that CFG++, initially designed to enhance text alignment in image diffusion models, also significantly improves image-to-video alignment in the video diffusion model. Specifically, after applying CFG++ into SVD sampling framework, the Euler-step of SVD (3) now reads:
where the last term in (3) is replaced by . In practice, we apply DDS guidance before CFG++ update, so in (14) should be replaced with as in (12), (13). We experimentally found that incorporating DDS and CFG++ guidance synergistically improves the interpolation performance of bidirectional sampling. The overall sampling method effectively steers the SVD sampling path to perform keyframe interpolation in an on-manifold manner, fully leveraging the generation capabilities of SVD. The detailed algorithm is provided in Algorithm 1. The vanilla bidirectional sampling can be implemented by removing DDS guidance (orange) and replacing the CFG++ update (blue) with a traditional CFG update. The detailed algorithm of the vanilla bidirectional sampling is provided in Appendix A.
@@ 4.2 Comparative studies
Quantitative evaluation. For quantitative evaluation, we used LPIPS (Zhang et al., 2018) and FID (Heusel et al., 2017) to assess the quality of the generated frames, and FVD (Unterthiner et al., 2019) to evaluate the overall quality of the generated videos. As shown in Table 1, our method surpasses the other baselines in terms of fidelity. Moreover, it achieves superior perceptual quality, particularly in scenarios involving dynamic motions (DAVIS), indicating that our approach effectively addresses the issue of deviations from the diffusion manifold, resulting in improved video generation quality. For a more comprehensive comparison, we present quantitative evaluations using PSNR and SSIM metrics in Appendix B.1. Nonetheless, it is important to note that these metrics primarily evaluate pixel-wise accuracy, which does not always align with human perceptual preferences. Within the scope of video generative models that align better with human perceptual preferences, our proposed method outperformed all baseline methods.
@@ 5 Conclusion
6 Acknowledgment
This work was partly supported by Institute for Information & communications Technology Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT)(RS-2019-II190075, Artificial Intelligence Graduate School Support Program(KAIST) and the National Research Foundation of Korea under Grant RS-2024-00336454.
@@ Appendix A Algorithm
Appendix B Additional results and discussion
B.1 Quantitative evaluation results
In addition to the perceptual quality evaluations presented in Table 1, we provide the PSNR and SSIM scores for our method and baseline models to offer a clearer understanding of our work. These metrics were obtained by comparing the generated video frames with their corresponding ground truth frames. As shown in Table 4, FILM achieved the highest PSNR and SSIM scores among the evaluated video generative models, which can be attributed to its training strategy that employs PSNR-oriented loss functions, such as the L1 loss calculated between the ground truth and output frames. However, within the scope of video generative models that align better with human perceptual preferences, our proposed method outperformed all baseline models, recording the highest PSNR and SSIM scores on both the DAVIS and Pexels datasets. These results highlight our model’s ability to generate videos that are both visually realistic and exhibit superior fidelity.
Method DAVIS Pexels PSNR SSIM PSNR SSIM FILM 22.5521 0.5346 31.3986 0.8418 TRF 16.4078 0.4040 19.2670 0.5346 Generative Inbetweening 16.0377 0.4255 20.8624 0.6334 Ours 17.5215 0.4387 21.4235 0.5939
B.2 Ablation studies on advanced guidances
To emphasize the importance of bidirectional sampling, we conducted additional ablation studies on the guidance techniques, CFG++ and DDS, using the Pexels dataset. These studies involved evaluating the results of the vanilla models and comparing them with the models enhanced by incorporating the guidance techniques.
As shown in Table 5, our vanilla bidirectional sampling consistently outperforms TRF across all metrics. This demonstrates that bidirectional sampling alone effectively addresses off-manifold issues, mitigating the artifacts commonly observed in TRF. Unlike Generative Inbetweening, which requires fine-tuning of the diffusion model, our method operates entirely in a training-free manner.
When comparing models enhanced with guidance techniques, our method demonstrates superior performance compared to both TRF and Generative Inbetweening. Notably, TRF also shows improvements with the addition of guidance techniques, as reflected in the quantitative metrics in Table 5 and the qualitative results available on our anonymous Project page. The results indicate more stable motion and better preservation of bounding frame information. However, its performance remains lower than that of our full method, highlighting the critical role of bidirectional sampling in addressing off-manifold issues and enabling the generation of more accurate and higher-quality videos. In addition, our proposed method exhibits a significantly higher degree of improvement compared to TRF. Since manifold guidance methods, such as CFG++, were originally designed for on-manifold sampling, our method can be synergistically enhanced with these guidance techniques for further improvement. In contrast, TRF faces challenges in achieving this synergistic improvement due to the generation of off-manifold samples, highlighting the differences in performance and effectiveness between the two approaches. In contrast, the integration of guidance techniques into Generative Inbetweening does not yield substantial improvements. This limited effect may be due to the interference between its fine-tuning process and rotated temporal attention, which could hinder the effectiveness of the guidance techniques.
We strongly encourage readers to visit our anonymous Website for a deeper understanding of the distinct roles of bidirectional sampling and guidance methods.
Method Vanilla CFG++ & DDS LPIPS FID FVD LPIPS FID FVD TRF 0.2222 80.618 880.97 0.2010 52.738 778.69 Generative Inbetweening 0.1523 40.470 746.26 0.1662 42.487 747.95 Ours 0.2074 63.241 717.37 0.1366 37.341 452.34
B.3 Discussions on the roles of CFG++ and DDS
CFG++. CFG++ is a guidance technique designed to improve alignment between images and text conditions. By applying CFG++ guidance to our model, the generated videos achieve better alignment with the image conditions and maintain semantic consistency with the given frames, , and . This enhancement positively impacts the perceptual quality of the videos.
DDS. SVD, an image-to-video diffusion model, generates 25-frame videos based on an initial frame condition, without utilizing a final frame condition. To address the absence of a last frame condition within the video interpolation framework, we incorporated DDS guidance. This approach ensures alignment of the last frame with in the temporally forward path and in the temporally backward path, thereby enhancing the temporal consistency of the generated videos.
B.4 Ablation study on SVD version.
Since we employed SVD-XT, which is the fine-tuned version of the original SVD, we checked the performance of our method with SVD. As shown in the Table 6, we see the performance increment in better video generation models.
Method (Pexels) FVD LPIPS SVD 608.07 0.1577 SVD-XT (ours) 452.34 0.1366
B.5 Future work and discussions
Our proposed method demonstrates exceptional performance in generating intermediate frames from two bounding frames. However, as the model is based on Stable Video Diffusion, which employs image embeddings as conditions for cross-attention instead of textual prompts, we have not considered the incorporation of text conditioning. Nonetheless, when bidirectional sampling is applied to other image-to-video (I2V) diffusion models, such as DynamiCrafter, we believe that it becomes feasible to guide actions based on textual conditions. Extending our method to support text-based control presents a promising avenue for future research.
B.6 Additional experimental results